OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 5x – 2y = 11? -11/2 -5/2 -2/5 5/2

OpenStudy (amoodarya):

5x – 2y = 11 first turn it to standard form of y=mx+h then m is slope 5x – 2y = 11 -2y=-5x+11 divide by -2

OpenStudy (amoodarya):

y=5/2 x-11/2 so slope is +5/2

OpenStudy (anonymous):

5/2